# Exploiting the Layered Intrinsic Dimensionality of Deep Models for Practical Adversarial Training

### v1

@@ 5.1 Robustifying Classification Models
2.4 ±0.1 41.4 ±0.2 4.0 ±0.2 2.4 ±0.0 2.4 ±0.0 5.6 ±0.4 2.4 ±0.1
Runtime efficiency. In Table 2, we provide details on the training time per epoch and inference time per instance for the IMDB dataset using BERT (RoBERTa has the same architecture). SMAAT has comparable efficiency to standard training and is on average, 3 times faster than standard AT during training. This is attributed to shorter backpropagation chains as AT in SMAAT is performed in the last layer ($l^{*}=l_{n}$). Note that FreeLB++, which inject noise in the first layer, remains inefficient even when PGD is replaced with FGSM [80, 32], as a full-depth backpropagation chain is still required. Certified defense baselines (SAFER, RSMI) and input space attacks (ASCC, DNE) are also time-consuming as they either require mapping samples into the manifold or performing an extensive search over word substitutes.
Figure 4: Effect of applying AT on generalization/robustness when applied to middle layers of BERT on the YELP dataset against TextFooler attack. Model robustness improves as adversarial training is applied to higher layers while generalization remains similar.
AT in intermediate layers. We experimentally show that, under decreasing ID behavior, AT in higher-layers yields increased robustness with increasing layer index. Thus, we apply AT to layers $l_{i}$ with $i\in\{0,2,4,6,8,10,11,\text{and }12\}$. We keep the number of PGD steps to 3 and the number of training epochs to 5 across layers. We report results on AT of BERT on the YELP dataset in Fig. 4. As expected, robustness (under TextFooler attack) increases with the layer index $i$ without affecting generalization. This observation aligns with our findings in Figure 3. The rise in OFM-AE ratio corresponds to an increase in robustness as we progress to higher layers.
Additional robustness evaluation of SMAAT on language understanding benchmarks, including GLUE and advGLUE are reported in Appendix D. Also, to emphasize that in the general case AT at the last layer is not sufficient for improving robustness, we conduct experiments as detailed in Appendix E.
5.2 Robustifying Safety Filters in Dec-LLMs
@@
Table 3: Robustifying safety filters based on dec-LLMs on AdvBench and HH_RLHF datasets. We report clean accuracy (CA), and robust accuracy under attack (AUA) which indicates the percentage of harmful prompts augmented with adversarial suffices accurately classified as harmful.
@@
Table 4: Robustifying retriever models of RAG on the Natural Questions dataset. We report Recall@10 (R@10) and Recall@100 (R@100) on the clean corpus for generalization. Robust recall (RR) is measured by how many samples are selected without adversarial passages in the top-k passages, with R@10 and R@100 corresponding to the top-10 and top-100 passages, respectively. During attacks, 10 and 50 adversarial passages are created, denoted as (N=10) and (N=50), respectively.
@@
The broad attack surface of LLMs has compelled model owners to implement solutions that extend beyond the safety alignment of a model. Notably, this includes content moderation filters that verify the harmlessness of a model’s inputs and outputs [27, 9]. These filters are essentially text classifiers, often based on lightweight enc-LLMs to minimize overhead and preserve the high responsiveness of dec-LLMs [27]. The robustness of safety filters is also a concern, as they too can be vulnerable to attacks. Adversarial attacks. The Greedy Coordinate Gradient (GCG) attack [81], which generates adversarial suffixes that are appended to the input prompts of a language model to bypass safety alignment, proves effective across both encoder-based and decoder-based architectures. Base models. We evaluate SMAAT’s effectiveness in improving the robustness of BERT, RoBERTa, and DistilBERT models when employed as safety filters. Dataset. Tests are performed on two datasets: the AdvBench dataset [81] and the Helpfulness-Harmfulness Human Preference dataset (HH-RLHF) [7]. We use GCG attack. Metrics. In our tests, we benchmarked the filter’s robustness by considering its generalization accuracy (ACC) for safe and harmful prompts, and robust accuracy under attack (AUA) by measuring the detection rate of harmful prompts augmented with adversarial suffixes. Further details are presented in Appendix F.
The results in Table 4 show that SMAAT significantly enhances the robustness of safety filters against the GCG attack across all models, while maintaining generalization accuracy compared to standard training. In line with our observation, FreeLB++ improves generalization as it applies AT in the first layer.
5.3 Robustifying Retriever Models of RAG
@@
Baselines. We evaluate the robustness of retriever models within the Retrieval-Augmented Generation (RAG) framework under standard, FreeLB++ and SMAAT. RAG combines a retriever model, which identifies relevant passages from a large corpus, with a generator model that constructs answers based on the retrieved information. Adversarial attacks. Poisoning attacks [78], which manipulate the retrieval corpus by generating adversarial passages, are particularly critical in this setup. Base model. We use the Contriever model [23], fine-tuned on the Natural Questions (NQ) [29] dataset, as our retriever. Metrics. We use robust recall (RR) measured by how many samples are selected without adversarial passages in the top-k passages, with R@10 and R@100 corresponding to the top-10 and top-100 passages, respectively.
Results in Table 4 show that the standard model offers limited robustness against the poisoning attack, with a significant number of adversarial passages being retrieved. FreeLB++ showed some improvement, reducing the number of adversarial passages retrieved. However, SMAAT significantly enhanced robustness, demonstrating a dramatic reduction in the retrieval of adversarial passages compared to the other methods. In terms of generalization, FreeLB++ yields the best results since it applies to AT in the initial layer (more ONM-AEs) while SMAAT and standard training have similar performance. In terms of generalization, FreeLB++ yields the best results as it applies AT in the first layer (resulting in more ONM-AEs), while both SMAAT and standard training exhibit similar performance.
@@ 6 Conclusion
In this paper, we introduce a manifold-aware approach designed to enhance the scalability, and robustness of adversarially trained deep neural networks while maintaining generalization. Our approach involves generating adversarial examples in higher layers with the lowest intrinsic dimensionality (ID). This strategy enables faster synthesis with shorter backpropagation chains and enhances robustness by capturing a higher proportion of off-manifold AEs. We achieve state-of-the-art performance in both runtime efficiency and robustness, while maintaining comparable levels of generalization to standard training. There are several avenues for future work including (i) investigate approaches for joint optimization of both generalization and robustness through perturbations in the intermediate layers, i.e., via controlling the on-off manifold AEs ratio (ii) study the impact of off-manifold AEs on out-of-distribution detection.
@@ Appendix A Supplementary Material
In this paper, we propose SMAAT, an adversarial training algorithm that not only optimizes for better robustness and generalization but also enhances scalability. Specifically, SMAAT leverages the manifold conjecture, which posits that off-manifold adversarial examples (OFM-AEs) lead to better robustness while on-manifold adversarial examples (ONM-AEs) enhance generalization. To achieve this, SMAAT perturbs the layer with the lowest intrinsic dimension (ID). Intuitively, such a layer is expected to yield the highest proportion of off-AEs. Formally, SMAAT solves the following program:
We show that $l^{*}$ corresponds to the last layer in encoder-based LLMs (enc-LLMs) and the first layer in decoder-based LLMs (dec-LLMs). This explains the difference in robustness and generalization trends between vision models and enc-LLMs [55, 48, 80]. Specifically, in vision models, improvements in robustness are often accompanied by a decrease in generalization. In contrast, for enc-LLMs, improvements in generalization are achieved with lesser gains in robustness.
SMAAT has the advantage of being significantly more runtime-efficient than classical adversarial training, by a factor $\mathcal{O}\Big{(}P(n-l^{*})\Big{(}\max\!\Big{(}\{d_{l_{i}}|l_{i}\in[1,l^{*}-1]\}\Big{)}-d_{l^{*}}\Big{)}\Big{)}$ for an $n$-layered deepnet under a $P$-step PGD attack where $d_{l}$ is the dimensionality of the layer $l$. Empirical results demonstrate that SMAAT leads to better robustness while maintaining comparable accuracy to standard training. Applications considered include robustifying (1) sentiment classifiers, (2) safety filters in decoder-based models, and (3) retriever models in the Retrieval Augmented Generation (RAG) setup.
The supplementary material is organized as follows:
@@
  • Appendix C: Additional Results of Robustifying Sentiment Classifiers
  • Appendix D: Additional Results of Language Understanding Benchmarks
  • Appendix E: Additional Results AT at the Last Layer
@@ Appendix B Intrinsic Dimension Estimation
In the following, we provide additional details on the experimental setup on off-manifold ratio calculation and ID estimation in Figure 3. Experimental setup. We conduct experiments on two enc-LLMs (BERT and RoBERTa), across three different datasets (AGNEWS, IMDB, YELP). Additionally, we evaluated three vision models (ViT [17], DeiT [58], VGG-13 [52]) on the CIFAR-10 [26] dataset and three LLMs (LLaMA-2-7b [59], Vicuna [77], Mistral-7b [24]) on the HelpSteer dataset  [66]. We utilize the train split of the datasets to estimate the ID and eigenspace of the layers. AEs are generated using TextFooler for LMs and PGD for vision models on the respective test splits of the datasets. For LLMs, AEs are generated using the Greedy Coordinate Gradient (GCG) attack [81] on the harmful prompt datasets released in the same paper. In our calculation, we use CLS embedding for BERT, RoBERTa, ViT, and DeiT models; last token embedding for LLMs; and conventional layer output for VGG-13.
@@
Appendix C Additional Results of Robustifying Sentiment Classifiers

### v2

@@ 5 SMAAT: Scalable Manifold Aware Adversarial Training
2.4 ±0.1 41.4 ±0.2 2.4 ±0.0 2.4 ±0.0 5.6 ±0.4 2.4 ±0.1
Complexity. Computing the layers IDs and searching for the optimal layer $l^{*}$ to perturb are done once per model and per task. Thus, they incur marginal overhead. For the AT part, when a $P$-step PGD attack is used, this results in $P$ forward-backward passes with length $(n-l^{*}+1)$ instead of $n$. The run-time of every forward/backward pass depends on the layer dimensionality, i.e., $O(d_{l-1}\times d_{l})$ for the $l^{\text{th}}$ layer. Overall, the complexity of one SMAAT forward-backward is $\mathcal{O}((n-l^{*}+1)\max_{l\in[l^{*},n]}(d_{l})^{2}))$. As a result, SMAAT is more efficient than standard AT by a factor of $\mathcal{O}(P\times l^{*}\cdot(\max_{i\in[1,n]}{(d_{i})^{2}}-\max_{j\in[l^{*},n]}{(d_{j})^{2}}))$. In the case Enc-LLMs (e.g., BERT and RoBERTa), $l^{*}$ is equal to $n$. The total run-time can be simplified to $\mathcal{O}(P\cdot(d_{n})^{2})$ where $d_{n}$ is the number of classes. Typical enc-LLMs tasks consist of less than five classes (Wang et al., 2019a) which makes the factor of classes small enough to be negligible. Hence, SMAAT enhances the efficiency of the AEs generation process by a factor of $l\cdot O(\max(d_{l}))$. This improvement practically eliminates the cost of the AE generation process.
6 Experiments
We assess robustbness, generalization and scalability of three types of models trained with SMAAT: (i) classifiers; (ii) retrievers in RAG settings and (iii) safeguarding models employed for moderating content produced by generative models. We consider four different attacks: (i) word substitution (Ren et al., 2019; Jin et al., 2020a; Li et al., 2020a), (ii) adaptive (Tramèr et al., 2018), (iii) Greedy Coordinate Gradient (GCG) (Zou et al., 2023), and (iv) corpus poisoning (Zhong et al., 2023) attacks. We include a run-time analysis to demonstrate scalability of SMAAT.
6.1 Sentiment and Topic Classifiers.
We adversarially trained two base models, BERT-base-cased and RoBERTa-base-cased on the tasks of sentiment classification (IMDB (Maas et al., 2011) and YELP (Zhang et al., 2015)) and topic classification (AGNEWS (Zhang et al., 2015)) under three input space attacks including PWWS (Ren et al., 2019) (synonym based), TextFooler (Jin et al., 2020a) (neighbor based), and BERT-Attack (Li et al., 2020a) (masked-LM based). Attacks are conducted using the TextAttack framework (Morris et al., 2020) and following the settings introduced by (Li et al., 2021b). We compare SMAAT to standard (non-adversarial) training, and six baselines from three families of defenses: (1) Input space AT (ASCC (Dong et al., 2021)), (2) Embedding space AT (FreeLB++ (Li et al., 2021b)), and (3) Certified defenses (SAFER (Ye et al., 2020), TMD (Minh & Luu, 2022), RSMI (Minh & Luu, 2022)). FreeLB++ focuses on scalability and robustness by minimizing the number of PGD steps and applying AT in the initial layer. TMD leverages manifold features by projecting samples back to the manifold in the last layer. Implementation details: In both base models, $l^{*}=n$, i.e., we generate AEs in the last layer. Specifically, we perturb the [CLS] embeddings before the classifier layer by freezing all layers before $l^{*}$. Further details are presented in Appendix D.
@@ 6 Experiments
Table 3: Robustifying dec-LLMs with encoder-based (BERT etc.) safety filters on AdvBench and HH_RLHF datasets. We report clean accuracy (CA), and robust accuracy under attack (AUA) which indicates the percentage of harmful prompts augmented with adversarial suffices accurately classified as harmful. Notice that SMAAT achieves very high robustness while other models sometimes completely fail.
@@
Robustness and generalization: In Table 1, we report results on robustness and generalization. On average, SMAAT demonstrates superior robustness over all datasets, with an improvement of 8.6%, 15.7%, and 28.8% over the best score for the BERT model, and 6.0%, 5.8%, and 19.0% for the RoBERTa model on the AGNEWS, IMDB, and YELP datasets, respectively. Note that FreeLB++, which perturbs the first layer, showed the best generalization in four out of six cases, as it produces mostly on-manifold examples (the first layer has a high dimension, as illustrated in Fig. 4) which is in line with the manifold conjecture. SMAAT maintains generalization in five out of six cases and only shows $0.5$ drop in performance in the case of RoBERTa with the IMDB dataset. Besides, SMAAT consistently outperforms TMD, even though it is a manifold-based method differently from SMAAT, TMD estimates the manifold and projects the input samples onto it before classification witch could lead to inaccuracies in the manifold estimation. SMAAT on the other hand is more robust to such inaccuracies as it only leverages the ID for selecting the layer to perturb. RSMI shows strong robustness against PWWS, as its masked inference mechanism inherently mimics synonym-based defenses without needing an explicit synonym set (Minh & Luu, 2022).
Runtime efficiency: In Table 2, we provide details on the training time per epoch and inference time per instance for the IMDB dataset using BERT (RoBERTa has the same architecture). SMAAT has comparable efficiency to standard training and is on average 3 times faster than standard AT during training. This is attributed to shorter backpropagation chains as AT in SMAAT is performed in the last layer ($l^{*}=n$). Note that FreeLB++, which injects noise in the first layer, remains inefficient even when PGD is replaced with FGSM (Zhu et al., 2020; Li et al., 2021b), as a full-depth backpropagation chain is still required. Certified defense baselines (SAFER, RSMI) and input space attack (ASCC) are also time-consuming as they either require mapping samples into the manifold or performing an extensive search over word substitutes. Additional evaluation of SMAAT on language understanding benchmarks, including GLUE and advGLUE are reported in Appendix E.
Table 4: Robustifying retriever models of RAG on the Natural Questions dataset. We report Recall@10 (R@10) and Recall@100 (R@100) on the clean corpus for generalization. Robust recall (RR) is measured by how many samples are selected without adversarial passages in the top-k passages, with R@10 and R@100 corresponding to the top-10 and top-100 passages, respectively. During attacks, 10 and 50 adversarial passages are created, denoted as (N=10) and (N=50), respectively. SMAAT makes RAG substantially more robust against selecting adversarial passages.
@@ 6.2 Safety Filters
The broad attack surface of LLMs has compelled model owners to implement solutions that extend beyond the safety alignment of a model. Notably, this includes content moderation filters that verify the harmlessness of a model’s inputs and outputs (Kumar et al., 2023; Cao et al., 2023; Fatehkia et al., 2025). These filters are essentially text classifiers, often based on lightweight enc-LLMs to minimize overhead and preserve the high responsiveness of dec-LLMs (Kumar et al., 2023). The robustness of safety filters is also a concern, as they too can be vulnerable to attacks. Base models: We evaluate SMAAT’s effectiveness on BERT, RoBERTa, and DistilBERT when used as safety filters. Datasets: The AdvBench dataset (Zou et al., 2023) and the Helpfulness-Harmfulness dataset (HH-RLHF) (Bai et al., 2022). Adversarial attack:, we use GCG attack. Metrics: we assess generalization accuracy (ACC) for safe and harmful prompts, and robust accuracy under attack (AUA) by measuring the detection rate of harmful prompts augmented with adversarial suffixes. Further details are presented in Appendix F. Robustness and generalization: Results in Table 3 show that SMAAT significantly enhances the robustness of safety filters against the GCG attack across all models, while maintaining generalization compared to standard training. Also, FreeLB++ improves generalization as it applies AT in the first layer.
6.3 Retriever Models of RAG
@@
RAG combines a retriever model, which identifies relevant passages from a large corpus, with a generator model that constructs answers based on the retrieved information. Base model: We use the Contriever model (Izacard et al., 2021), fine-tuned on the Natural Questions (NQ) (Kwiatkowski et al., 2019) dataset, as our retriever. Baselines: We evaluate the robustness of retriever models within the RAG framework under standard, FreeLB++ and SMAAT. Adversarial attacks:, we use poisoning attacks (Zhong et al., 2023), which manipulate the retrieval corpus by generating adversarial passages. As metrics, we use robust recall (RR) measured by how many samples are selected without adversarial passages in the top-k passages, with R@10 and R@100 corresponding to the top-10 and top-100 passages, respectively. Robustness and generalization: Table 4 shows that the standard model offers limited robustness against the attack, with a significant number of adversarial passages being retrieved. FreeLB++ shows some improvement, reducing the number of adversarial passages retrieved. However, SMAAT significantly enhances robustness, demonstrating a dramatic reduction in the retrieval of adversarial passages compared to the others. In terms of generalization, FreeLB++ yields the best results as it applies AT in the first layer (resulting in more ONM-AEs), while both SMAAT and standard training exhibit similar performance.
@@ 7 Conclusion
In this paper, we explain the varying effects of AT across architectures using the manifold conjecture, showing that ID influences robustness and generalization. Vision and decoder-based models favor off-manifold AEs in early layers, enhancing robustness but harming generalization, while encoder-based LLMs favor on-manifold AEs, preserving generalization with limited robustness gains. Leveraging this insight, we introduce SMAAT, which improves AT scalability by perturbing the layer with the lowest ID. This reduces training overhead by 25–33% while boosting robustness across sentiment classification, safety filtering, and retrieval tasks, maintaining generalization. We plan to explore joint optimization of generalization and robustness by controlling the on/off-manifold AE ratio through perturbations in intermediate layers.
Acknowledgement
This work is partial supported by the Qatar National Research Fund (QNRF) grant NPRP11C-1229-170007.
Impact Statement
This research aims to deepen the understanding of how AT methods influence generalization and robustness, leading to more effective and practical adversarial training techniques. As machine learning models become increasingly integrated into real-world applications, it is crucial to develop frameworks that mitigate misuse while ensuring ethical compliance. Building on prior work in adversarial training, our approach enhances the robustness of encoder-based models while significantly improving computational efficiency.
@@ Appendix A Supplementary Material
In this paper, we propose SMAAT, an AT algorithm that not only optimizes for better robustness and generalization but also enhances scalability. Specifically, SMAAT leverages the manifold conjecture, which posits that OFM-AEs lead to better robustness while ONM-AEs enhance generalization. To achieve this, SMAAT perturbs the layer with the lowest intrinsic dimension (ID). Intuitively, perturbing this layer would yield the highest proportion of OFM-AEs across layers. Formally, SMAAT solves the following program:
$\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big{[}\max_{\|\delta_{l^{*}}\|\leq\epsilon_{l^{*}}}\ell\Big{(}f_{\theta^{[l^{*},n]}}\Big{(}f_{\theta^{[1,l^{*}]}}(x)+\delta_{l^{*}}\Big{)},y\Big{)}\Big{]}.$
We show that $l^{*}$ corresponds to the last layer in enc-LLMs and the first layer in dec-LLMs. This explains the difference in robustness and generalization trends between vision models and enc-LLMs (Stutz et al., 2019a; Shafahi et al., 2019; Zhu et al., 2020). Specifically, in vision models, improvements in robustness are often accompanied by a decrease in generalization. In contrast, in enc-LLMs, improvements in generalization are achieved with lesser gains in robustness. The algorithm for SMAAT is provided in Alg. 1.
Algorithm 1 SMAAT
1:  Input: $\mathcal{D}=\{\mathcal{\bm{X}},\mathcal{\bm{Y}}\}$: input data, $f_{\theta}$: a deepnet model, $\epsilon$: attack strength, $E$: the number of epochs, $\alpha$: PGD learning rate, and $\Pi$: the projection operator into the $\epsilon$-ball.
2:  Output $f_{\theta}$: a deepnet model
3:   % Determine the ID behaviour of the model with twoNN
4:   % Identify the optimal layer $l^{*}$ to perturb (Eq. (7))
5:  for $e=1,..,E$ do
6:     for $(x_{i},y_{i})\in\mathcal{D}$ do
7:        $\delta_{l^{*}}\sim\mathcal{N}(0,\sigma^{2}I)$ sample an initial perturbation
8:         % store forward pass for scalability
9:        $mid\_rep\leftarrow f_{\theta^{[0,l^{*}]}}(x_{i})$
10:        for $s=1,..,S$ do
11:           $loss\leftarrow\ell\left(f_{\theta^{[l^{*},n]}}(mid\_rep+\delta_{l^{*}}),y_{i})\right))$
12:           $\delta_{l^{*}}\leftarrow\Pi_{\epsilon_{l^{*}}}(\delta_{l^{*}}+\alpha\cdot sign\left(\nabla_{\delta_{l^{*}}}(loss))\right)$
13:        end for
14:         % update the model parameters
15:        $loss\leftarrow\ell\left(f_{\theta^{[l^{*},n]}}(mid\_rep+\delta_{l^{*}}),y_{i})\right)$
16:        $\theta^{[l^{*},n]}\leftarrow\theta^{[l^{*},n]}-lr\nabla_{\theta^{[l^{*},n]}}(loss)$
17:     end for
18:  end for
SMAAT has the advantage of being significantly more runtime-efficient than classical adversarial training. The gain is of the order of $\mathcal{O}\Big{(}P(n-l^{*})\Big{(}\max\!\Big{(}\{d_{l_{i}}|l_{i}\in[1,l^{*}-1]\}\Big{)}-d_{l^{*}}\Big{)}\Big{)}$, for an $n$-layered deepnet under a $P$-step PGD attack. $d_{l}$ is the dimensionality of the layer $l$.
Empirical results demonstrate that SMAAT leads to better robustness while maintaining comparable accuracy to standard training. We achieve compelling results on robustifying (1) sentiment classifiers, (2) safety filters in decoder-based models, and (3) retriever models in the Retrieval Augmented Generation (RAG) setup.
The remaining of the supplementary material is organized as follows:
@@
  • Appendix C: Additional Results of ID & Robustness/Generalization Relationship
  • Appendix D: Additional Results of Robustifying Sentiment Classifiers
  • Appendix E: Additional Results of Language Understanding Benchmarks
@@ Appendix B Intrinsic Dimension Estimation
In the following, we provide additional details on the experimental setup on OFM-/ONM- AEs ratio calculation and the ID estimation in Fig. 4.
Experimental setup. We conduct experiments on two enc-LLMs (BERT and RoBERTa), across three different datasets (AGNEWS, IMDB, YELP). Additionally, we evaluated three vision models (ViT (Dosovitskiy et al., 2020), DeiT (Touvron et al., 2021), VGG-13 (Simonyan & Zisserman, 2014)) on the CIFAR-10 (Krizhevsky, 2009) dataset and three LLMs (LLaMA-2-7b (Touvron et al., 2023), Vicuna (Zheng et al., 2024), Mistral-7b (Jiang et al., 2023)) on the HelpSteer dataset  (Wang et al., 2023). We utilize the train split of the datasets to estimate the ID and eigenspace of the layers. AEs are generated using TextFooler for LMs and PGD for vision models on the respective test splits of the datasets. For LLMs, AEs are generated using the Greedy Coordinate Gradient (GCG) attack (Zou et al., 2023) on the harmful prompt datasets released in the same paper. In our calculation, we use CLS embedding for BERT, RoBERTa, ViT, and DeiT models; last token embedding for LLMs; and conventional layer output for VGG-13.
@@
Appendix C Additional Results of ID & Robustness/Generalization Relationship
We further examine the relationship between intrinsic dimensionality (ID), robustness, and generalization by applying Latent Adversarial Training (LAT) to different layers of the LLaMA-2-7B model. Unlike untargeted LAT, which disrupts model behavior, we employ targeted LAT to induce specific adversarial responses, following Casper et al. (2024). This is achieved by perturbing the residual stream with L2-norm-bounded noise using PGD.
To stabilize training and mitigate unintended effects, we interleave LAT with supervised fine-tuning on the UltraChat dataset (Ding et al., 2023). We evaluate generalization using MT-Bench (Zheng et al., 2023) and robustness using GCG-based and PAIR attacks from HarmBench (Mazeika et al., 2024). The model is trained with a learning rate of $42e-$ 4, applying LAT at every even-numbered layer with norm bounds ranging from 1 to 5.
The results are presented in Figure 6. Attacking lower layers (light blue markers) improves robustness but reduces generalization, as indicated by points concentrated in the bottom-right region. In contrast, attacking upper layers (dark blue markers) enhances generalization at the cost of robustness, shifting results toward the top-left. Additionally, while model robustness drops to as low as 30% under the GCG attack, it remains consistently above 75% across all configurations under the PAIR attack.
Appendix D Additional Results of Robustifying Sentiment Classifiers
(a) GCG
(b) PAIR
Figure 6: Layer-wise effects of adversarial training on generalization and robustness against GCG and PAIR attacks on the LLaMA-2 model. Marker colors represent layer depth, transitioning from light blue (lower layers) to dark blue (higher layers). The results show that robustness drops to as low as 30% under the GCG attack, while remaining above 75% under the PAIR attack across all configurations.